# On measuring grounding and generalizing grounding problems

### v1

@@ 8.2 A typology of grounding profiles
Table 4: Fourfold typology for robustness and compositionality.
@@
Table 5: Grouding archetypes, characterized by their respective typological behaviors.
@@ Appendix B Toy example
To demonstrate the application of the framework and clarify the independence of the desiderata, we detail here an extremely simplified toy example, a concrete evaluation of a toy agent in grid-world, of which we implemented and trained via policy gradient reinforcement learning. Code for this example may be found at the project repo.
@@ B.1 Evaluation setup
We define the evaluation tuple $E=(k,t,U,P)$. The context $k$ is a continuous $10\times 10$ planar world; meaning type $t$ is extensional, where meanings correspond to 2D coordinates; threat model $U$ Gaussian noise injected into the representation space $\mathcal{R}$; reference distribution $P$ uniform over valid color-direction command pairs.
The agent’s vocabulary consists of atoms $\Sigma_{\text{atom}}=\{\text{RED},\text{BLUE},\text{NORTH},\text{SOUTH},\text{EAST},\text{WEST}\}$. The intended interpretation $\mathcal{I}_{k}^{t}$ maps colors to landmarks and directions to unit vectors:
@@
The semantic algebra $\langle\mathcal{M}_{k}^{t},\mathcal{F}_{k}^{t}\rangle$ interprets the grammar’s composition constructor as vector addition:
@@ B.2 Grounding architecture
  • $\Phi:\Sigma^{*}\to\mathcal{R}$ embedding layer followed by a GRU, mapping token sequences to hidden states; the representation space is $\mathcal{R}=\mathbb{R}^{64}$ with Euclidean metric $d_{\mathcal{R}}$;
  • $\Gamma:\mathcal{R}\to\mathcal{C}$ linear decoder mapping hidden states to predicted coordinates; concept space is $\mathcal{C}=\mathbb{R}^{2}$;
@@
The agent was trained using the classic REINFORCE algorithm over 3000 episodes, minimizing Euclidean distance to target coordinates. Training logs (an excerpt of which is shown in Table 6) confirm convergence from an initial loss of $5.23$ to near-zero:
Table 6: Excerpt of training logs.
episode 0: loss 5.234 (untrained)
episode 500: loss 0.0002
episode 2500: loss 0.0033 (converged)
@@ B.3 Grounding audit
Let us now walk through the grounding desiderata.
@@
The mapping mechanisms ($\Phi$, $\Gamma$) are neural network weights acquired via reward-driven optimization; the system satisfies G0-strong: the semantic mappings are both implemented internally and acquired through a selection process ($\mathcal{T}$ RL training) relevant to the task family.
@@
$\mathcal{A}_{k}^{t}\Psi(\text{RED})=(8.027,7.987),\quad\mathcal{I}_{k}^{t}(\text{RED})=(8.0,8.0).$
@@
$\varepsilon_{\text{pres}}^{k,t}=\|(8.027,7.987)-(8.0,8.0)\|_{2}=0.030.$
On a $10\times 10$ world, this represents drift of $0.3\%$ of the total scale.
@@
$\mathcal{A}_{k}^{t}\Psi(\text{RED NORTH})=(8.041,9.003),\quad\mathcal{I}_{k}^{t}(\text{RED NORTH})=(8.0,9.0).$
@@
$\varepsilon_{\text{faith}}^{k,t}=\|(8.041,9.003)-(8.0,9.0)\|_{2}=0.041.$
This is modestly larger than the atomic preservation error: accumulated drift across the two constituents.
G2b (etiological faithfulness)
@@
  • $M$ active: output $(8.041,9.003)$, task success (distance <0.5<0.5).
  • $M$ ablated: output $(8.027,7.987)$, task failure.
The ablated output equals $\mathcal{A}_{k}^{t}\Psi(\text{RED})$ exactly: the agent navigates to the landmark, indeed, but ignores the directional modifier entirely. The average causal effect is $\text{ACE}_{E}(M)=1.0$, the modifier-processing mechanism is strictly necessary for composed commands.
@@
  • Unperturbed output: $(8.027,7.987)$;
  • Perturbed output: $(8.060,7.905)$;
  • Semantic drift: $0.088$.
The local modulus satisfies $\omega_{U}^{k,t}(0.5)=0.088$, yielding a ratio of approximately $0.18$: less than one-fifth of the representational perturbation propagates to semantic output. This dampening follows from the dimensionality reduction through $\Gamma$ ($\mathbb{R}^{64}\to\mathbb{R}^{2}$) and the learned smoothness of the decoder. The behavior contrasts sharply with the degenerate step-function modulus of uniformly discrete symbolic systems (Corollary A.1), where any perturbation exceeding the minimal symbol spacing can induce arbitrary semantic change.
@@
$\mathcal{A}_{k}^{t}\Psi(\text{RED})$ $=(8.027,7.987),$
$\mathcal{A}_{k}^{t}\Psi(\text{NORTH})$ $=(-0.009,0.994).$
The agent’s learned representation of NORTH is nearly the true unit vector $(0,1)$, with a small spurious negative $x$-component. If the agent were perfectly compositional:
$f_{k}^{\mathcal{M},t}\big(\mathcal{A}_{k}^{t}\Psi(\text{RED}),\,\mathcal{A}_{k}^{t}\Psi(\text{NORTH})\big)=(8.027,7.987)+(-0.009,0.994)=(8.018,8.981).$
The actual output was $(8.041,9.003)$. The compositionality deficit is:
$\delta_{\text{comp}}^{k,t}=\|(8.041,9.003)-(8.018,8.981)\|_{2}=0.031.$
@@ B.4 Summary remarks
Taken together, we observe the following desiderata and values in Table 7.
Table 7: Grounding profile for gridworld agent.
@@
G1 (preservation) $\varepsilon_{\text{pres}}^{k,t}=0.030$ atoms stable
G2a (faithfulness) $\varepsilon_{\text{faith}}^{k,t}=0.041$ composed outputs accurate
G2b (etiological) $\text{ACE}_{E}(M)=1.0$ strictly necessary mechanism
G3 (robustness) $\omega_{U}^{k,t}(0.5)=0.088$ strong dampening
G4 (compositionality) $\delta_{\text{comp}}^{k,t}=0.031$ highly systematic
First, $\varepsilon_{\text{faith}}^{k,t}>\varepsilon_{\text{pres}}^{k,t}$ (0.041 versus 0.030): the composed expression incurs modestly more error than the atoms alone; this is expected when composition accumulates independent atomic errors, though the increase here is slight.
Second, $\delta_{\text{comp}}^{k,t}\approx\varepsilon_{\text{pres}}^{k,t}$ (0.031 versus0.030): the compositionality deficit is comparable to the atomic preservation error; the agent’s deviation from perfect composition is attributable primarily to imperfect atoms rather than to failures of the combinatory mechanism itself. The agent has learned to add correctly; it merely adds slightly wrong vectors.
Third, the robustness ratio $\omega_{U}^{k,t}(0.5)/0.5\approx 0.18$ quantifies the degree to which the decoder $\Gamma$ compresses high-dimensional representational noise into low-dimensional semantic output. This compression is a structural property of the architecture (projecting $\mathbb{R}^{64}$ onto $\mathbb{R}^{2}$) combined with learned smoothness; a symbolic system with discrete $\mathcal{R}$ would exhibit no such compression.
Finally, the ablation study (G2b) confirms that the measured success is not accidental: the sequential processing of modifiers is strictly necessary, and removing it reduces the agent to landmark-only navigation. The ACE of $1.0$ establishes maximal etiological warrant for the modifier-integration mechanism.
Our grid-world shows that G2a and G4 are orthogonal, in that an agent could exhibit perfect compositionality ($\delta_{\text{comp}}^{k,t}=0$) by consistently applying addition to incorrect atoms, even if both the parts and the whole were far from the ground truth; conversely, an agent could achieve perfect faithfulness on a phrase by memorizing the correct output without systematic combination of parts (high $\delta_{\text{comp}}^{k,t}$).
The agent demonstrates a different pattern: moderate accuracy with high systematicity. Its errors are primarily calibration errors in the atomic concepts. The grounding profile replaces the question “Does the agent understand?” with a quantified answer: it preserves atoms within $0.030$ units, composes them with $0.031$ units of systematic deviation, dampens representational noise by a factor of five, and its semantic mechanisms are causally necessary for task success. Whether this profile suffices depends on the application, of course: an assistant requiring $0.01$-unit precision would reject this agent; an assistant tolerating $0.1$-unit would accept it. The framework provides the measurements; adequacy thresholds remain domain-specific.

### v2

@@ 1 Introduction
A pictorial representation is given in Figure 1, upon which we will later populate with our formalisms proper (Figure 2) and an example evaluation (Figure 3): an evaluation tuple parametrizes the audit of a grounding architecture, which processes symbols through representations and concepts to meanings, compared against an intended interpretation, the instantiations of which give particular modes. Desiderata (authenticity, preservation, faithfulness, robustness, compositionality) yield a grounding profile that locates the system within a typology of archetypes.
[…]
Figure 1: An evaluation tuple fixes the context (e.g., logical, embodied, linguistic, or multimodal settings), the meaning type (extensional, inferential, or social-normative), the threat model (what perturbations the system must withstand), and the reference distribution (against which robustness is measured). The grounding architecture comprises a processing chain: surface symbols are encoded into internal representations, which are then organized into task-level concepts, and finally aligned to a space of meanings relevant to the evaluation. The intended interpretation provides the gold-standard meanings against which the agent’s outputs are compared. The architecture may instantiate different modes: symbolic (meaning via rules); referential (meaning via world-contact); vectorial (meaning via geometry); relational (meaning via inferential role). Desiderata audit the architecture: (G0) authenticity asks whether semantic mechanisms are internal and earned; (G1) preservation asks whether atomic meanings survive processing; (G2) faithfulness asks whether outputs match targets and whether this match was selected-for; (G3) robustness asks whether meanings degrade gracefully under noise; (G4) compositionality asks whether wholes derive systematically from parts. The audit produces a grounding profile, a set of measured errors and parameters that determines the system’s typology: whether it behaves as a “grounded” system, a “parrot” (accurate but unsystematic), a “brittle” expert (accurate but fragile), or any of other archetypes.
@@ 3.1 Formalization
The same structure as in Figure 1 is given in Figure 2 with formal notation: surface symbols $\Sigma$ map via $\Phi$ to representations $\mathcal{R}$, via $\Gamma$ to concepts $\mathcal{C}$, and via alignment $\mathcal{A}_{k}^{t}$ to meanings $\mathcal{M}_{k}^{t}$. The evaluation tuple $E=(k,t,U,P)$ indexes all measurements. The grounding profile $\text{GP}(\mathfrak{G};E)$ collects preservation error $\varepsilon_{\text{pres}}^{k,t}$, faithfulness error $\varepsilon_{\text{faith}}^{k,t}$, causal effect $\text{ACE}_{E}(M)$, robustness modulus $\omega_{U}^{k,t}$, composition error $\delta_{\text{comp}}^{k,t}$, and systematicity $\beta^{k,t}$.
[…]
Figure 2: An evaluation tuple $E=(k,t,U,P)$ fixes the context, meaning type, threat model, and reference distribution, parametrizing the audit of a grounding architecture $\mathfrak{G}$. Within $\mathfrak{G}$, the processing chain maps surface symbols $\Sigma$ through an encoder $\Phi$ to internal representations $\mathcal{R}$, then via $\Gamma$ to task-level concepts $\mathcal{C}$, and finally through the alignment $\mathcal{A}_{k}^{t}$ to the meaning space $\mathcal{M}_{k}^{t}$ appropriate to the declared $(k,t)$. The architecture may instantiate different modes (symbolic, referential, vectorial, or relational) based on structural and etiological constraints. Desiderata (G0–G4) audit the architecture: authenticity; preservation; faithfulness; robustness; compositionality. The audit yields a grounding profile $\text{GP}(\mathfrak{G};E)=\{\varepsilon_{\text{pres}}^{k,t},\varepsilon_{\text{faith}}^{k,t},\text{ACE}_{E}(M),\omega_{U}^{k,t}(\cdot),\delta_{\text{comp}}^{k,t},\beta^{k,t}\}$ of measured parameters that locates the system within a typology of grounding archetypes.
@@ 8.2 A typology of grounding profiles
Table 4: Fourfold typology for robustness and compositionality
@@
Table 5: Grouding archetypes, characterized by their respective typological behaviors
@@ Appendix B Toy example
Let us demonstrate the application of the framework, and clarify the independence of the desiderata in a toy example of an agent in grid-world, of which we implemented and trained via policy gradient reinforcement learning. Code for this example may be found at the project repo.
@@ B.1 Evaluation setup
We define the evaluation tuple $E=(k,t,U,P)$. The context $k$ is a continuous $10\times 10$ planar world; meaning type $t$ is extensional, where meanings correspond to 2D coordinates; threat model $U$ Gaussian noise injected into the representation space $\mathcal{R}$; reference distribution $P$ uniform over valid color-direction command pairs. The agent’s vocabulary consists of atoms $\Sigma_{\text{atom}}=\{\text{RED},\text{BLUE},\text{NORTH},\text{SOUTH},\text{EAST},\text{WEST}\}$. The intended interpretation $\mathcal{I}_{k}^{t}$ maps colors to landmarks and directions to unit vectors:
@@
The algebra $\langle\mathcal{M}_{k}^{t},\mathcal{F}_{k}^{t}\rangle$ interprets the grammar’s composition constructor as vector addition:
@@ B.2 Grounding architecture
  • $\Phi:\Sigma^{*}\to\mathcal{R}$ is an embedding layer followed by a GRU, mapping token sequences to hidden states; the representation space is $\mathcal{R}=\mathbb{R}^{64}$ with Euclidean metric $d_{\mathcal{R}}$;
  • $\Gamma:\mathcal{R}\to\mathcal{C}$ a linear decoder mapping hidden states to predicted coordinates; the concept space is $\mathcal{C}=\mathbb{R}^{2}$;
@@
The agent was trained using the classic REINFORCE algorithm over 3,000 episodes, minimizing Euclidean distance to target coordinates. Two combinations (BLUE EAST and RED WEST) were held out during training to test systematicity. Training logs (an excerpt of which is shown in Table 6) confirm convergence from an initial loss of $3.21$ to near-zero:
Table 6: Excerpt of training logs
Episode loss
0 3.205 (untrained)
500 0.009
2500 0.0001 (converged)
@@ B.3 Grounding audit
Let us now walk through the grounding.
@@
The mappings ($\Phi$, $\Gamma$) are neural network weights acquired via reward-driven optimization. The system satisfies G0-strong: the semantic mappings are both implemented internally and acquired through a selection process ($\mathcal{T}$ RL training) relevant to the task family.
@@
$\mathcal{A}_{k}^{t}\Psi(\text{RED})=(7.941,8.224),\quad\mathcal{I}_{k}^{t}(\text{RED})=(8.0,8.0).$
@@
$\varepsilon_{\text{pres}}^{k,t}=\|(7.941,8.224)-(8.0,8.0)\|_{2}=0.231.$
On a $10\times 10$ world, this represents drift of the total scale, an otherwise notable miscalibration of the atomic concept.
@@
$\mathcal{A}_{k}^{t}\Psi(\text{RED NORTH})=(7.726,9.522),\quad\mathcal{I}_{k}^{t}(\text{RED NORTH})=(8.0,9.0).$
@@
$\varepsilon_{\text{faith}}^{k,t}=\|(7.726,9.522)-(8.0,9.0)\|_{2}=0.590.$
This substantially exceeds both the atomic preservation error and the task success threshold of $0.5$ units; in this way, the agent arrives in the correct region, but misses the target.
G2b (etiological faithfulness).
@@
  • $M$ active gives output $(7.726,9.522)$ and task failure (distance $\geq 0.5$);
  • $M$ ablated gives output $(7.941,8.224)$, and task failure.
The ablated output equals $\mathcal{A}_{k}^{t}\Psi(\text{RED})$ exactly: the agent navigates to its (miscalibrated) landmark concept, but ignores the directional modifier entirely! The average causal effect is $\text{ACE}_{E}(M)=0.0$, since both conditions fail the binary success criterion.
This result is notworthy, and requires careful interpretation here. Indeed, the mechanism is causally active: it shifts the $y$-coordinate from $8.224$ (ablated) to $9.522$ (active), a movement of about $1.3$ units in the correct direction; however, because neither condition crosses the success threshold, the binary ACE registers zero. This is a limitation of threshold-based success predicates: a mechanism can be genuinely load-bearing and yet fail to register when baseline performance is poor. Note: a continuous formulation of ACE (measuring distance reduction rather than success-rate difference) we expect to reveal the positive contribution.
@@
  • unperturbed output: $(7.941,8.224)$;
  • perturbed output: $(8.036,8.372)$;
  • semantic drift: $0.176$.
The local modulus satisfies $\omega_{U}^{k,t}(0.5)=0.176$, yielding a ratio of approximately $0.35$: roughly one-third of the representational perturbation propagates to semantic output. This dampening follows from the dimensionality reduction through $\Gamma$ (that is, $\mathbb{R}^{64}\to\mathbb{R}^{2}$), and is a structural property of the architecture, independent of the agent’s accuracy. Even a miscalibrated agent benefits from this compression. The behavior contrasts sharply with the degenerate step-function modulus of uniformly discrete symbolic systems (Corollary A.1), where any perturbation exceeding the minimal symbol spacing can induce arbitrary semantic change.
@@
$\mathcal{A}_{k}^{t}\Psi(\text{RED})$ $=(7.941,8.224),$
$\mathcal{A}_{k}^{t}\Psi(\text{NORTH})$ $=(-0.097,1.114).$
The agent’s learned representation of NORTH overshoots the true unit vector $(0,1)$ in the $y$-component and includes a spurious negative $x$-component. If the agent were perfectly compositional:
$f_{k}^{\mathcal{M},t}\big(\mathcal{A}_{k}^{t}\Psi(\text{RED}),\,\mathcal{A}_{k}^{t}\Psi(\text{NORTH})\big)=(7.941,8.224)+(-0.097,1.114)=(7.844,9.338).$
The actual output was $(7.726,9.522)$. The compositionality deficit is:
$\delta_{\text{comp}}^{k,t}=\|(7.726,9.522)-(7.844,9.338)\|_{2}=0.219.$
Unlike a well-calibrated agent where $\delta_{\text{comp}}^{k,t}\approx\varepsilon_{\text{pres}}^{k,t}$ would indicate correct algebra applied to slightly wrong atoms, here both values are large: the agent learned neither precise atoms nor perfect addition.
G4 (systematicity)
To assess whether the agent’s compositional behavior generalizes, we test on two combinations held out during training:
Command Output Error Success
BLUE EAST $(2.609,1.793)$ $0.442$ successful
RED WEST $(7.050,8.680)$ $0.682$ unsuccessful
The systematicity score is $\beta^{k,t}=0.50$: the agent generalizes to some novel combinations but not others. The asymmetry is notable, in that BLUE compositions succeed while RED compositions fail, which we account for as the underlying miscalibration of the RED landmark observed in G1: errors in atomic representations propagate to novel compositions, even when the compositional mechanism itself is intact.
@@ B.4 Summary remarks
Taken together, we observe the grounding profile shown in Table 7.
Table 7: Grounding profile for grid-world agent.
@@
G1 (preservation) $\varepsilon_{\text{pres}}^{k,t}=0.231$ atoms miscalibrated
G2a (faithfulness) $\varepsilon_{\text{faith}}^{k,t}=0.590$ composed outputs inaccurate
G2b (etiological) $\text{ACE}_{E}(M)=0.0$ mechanism active, below threshold
G3 (robustness) $\omega_{U}^{k,t}(0.5)=0.176$ moderate dampening
G4 (compositionality) $\delta_{\text{comp}}^{k,t}=0.219$ imperfect systematicity
G4 (systematicity) $\beta^{k,t}=0.50$ partial generalization
This agent is diagnostically interesting precisely because it is imperfect; the profile shows us a specific failure mode for which we qualified in our typology. Following Section 8.2, the agent is miscalibrated cell (low G2a, moderate G4): it has learned some compositional structure but applies it to poorly calibrated atoms; if we consider G2a $\times$ G2b, it represents an effortful failure, in that it has genuinely learned mechanisms (G0-strong, and the modifier mechanism does change output substantially), but those mechanisms do not achieve task success.
We now decorate our Figures 1 and 2 with our toy example, shown in Figure 3.
[…]
Figure 3: The evaluation tuple $E=(k,t,U,P)$ specifies: context $k$ as planar navigation; meaning type $t$ as extensional (2D coordinates); threat model $U$ as Gaussian noise injected into the representation space; reference distribution $P$ as uniform over color-direction command pairs. The grounding architecture implements: an embedding layer followed by a GRU encoder ($\Phi:\Sigma^{*}\to\mathbb{R}^{64}$), a linear decoder ($\Gamma:\mathbb{R}^{64}\to\mathbb{R}^{2}$), and an identity alignment ($\mathcal{A}_{k}^{t}$) to the meaning space. The intended interpretation $\mathcal{I}_{k}^{t}$ maps color atoms to landmark coordinates (e.g., RED $\mapsto(8,8)$) and direction atoms to unit vectors (e.g., NORTH $\mapsto(0,1)$), with composition defined as vector addition. The symbol-to-concept map $\Psi=\Gamma\circ\Phi$ compresses 64-dimensional hidden states to 2D coordinates, while the observable semantics $S_{k,t}=\mathcal{A}_{k}^{t}\circ\Gamma$ (here, trivially, since $\mathcal{A}_{k}^{t}$ is identity) delivers those coordinates as world positions. The dimensionality reduction through $\Gamma$ is responsible for the robustness compression observed in G3. This isntantiation qualifies the grounding as vectorial. The audit yields the profile: preservation error $\varepsilon_{\text{pres}}^{k,t}=0.2313$; faithfulness error $\varepsilon_{\text{faith}}^{k,t}=0.5897$; average causal effect $\text{ACE}_{E}(M)=0.0$; robustness modulus $\omega_{U}^{k,t}(0.5)=0.1760$; compositionality deficit $\delta_{\text{comp}}^{k,t}=0.2191$; systamaticity $\beta^{(k,t)}=0.5$. The profile classifies the agent as miscalibrated and an effortful failure.
Several observations clarify the failure.
  1. 1.
    The root cause is G1. The atomic concept of RED drifts by $0.231$ units from its target, and this miscalibration propagates through composition: even if the agent added perfectly, it would arrive at the wrong location.
  2. 2.
    The modifier-processing mechanism shifts output by $1.3$ units in the $y$-direction, which is exactly the kind of causal contribution G2b is designed to detect! The zero ACE reflects the binary success predicate, in that both conditions fail, so the difference is zero. Continuous, or, at least, margin-based formulations of etiological faithfulness, may better capture causal contribution when baseline performance is poor.
  3. 3.
    The robustness ratio $\omega_{U}^{k,t}(0.5)/0.5\approx 0.35$ demonstrates that architectural properties (dimensionality reduction, learned smoothness) provide robustness independent of accuracy (!). A miscalibrated agent still benefits from the compression of $\mathbb{R}^{64}\to\mathbb{R}^{2}$; robustness is not a reward for getting things right.
  4. 4.
    The asymmetry in $\beta^{k,t}$ (BLUE EAST succeeds; RED WEST fails) traces directly to the asymmetry in atomic calibration: the BLUE landmark is better learned than RED, so compositions involving BLUE generalize while those involving RED do not. Systematicity, we should note, is not a monolithic property, and it can vary across the vocabulary depending on which atoms are well-grounded.
The framework thus tells us that the agent fails and how it does so: miscalibrated atoms (G1) cascade to poor faithfulness (G2a) and partial systematicity ($\beta$), while the compositionality (G4) and robustness (G3) remain partially intact. A specific intervention follows: improve G1 through longer training, curriculum emphasizing atomic commands, or targeted data augmentation for RED combinations; a binary “grounded/ungrounded” judgment would miss this structure entirely.
Now, whether this profile is acceptable or unacceptable depends on the application: an agent tolerating $0.7$-unit error would pass; one requiring $0.1$-unit precision would fail. The framework simply provides the language for measurements and localizes the deficits; adequacy thresholds remain task- and domain-specific.